# Continuous-time q-learning in jump-diffusion models under Tsallis entropy

### v3

@@ 2.3 Exploratory HJB equation and policy improvement iteration
For a KKT multiplier $\xi\in\mathds{R}$, we consider the following unconstrained problem:
$\sup_{\pi\in{\cal P}(U)}\left[\int_{U}\left(q(u)\pi(u)-\frac{1}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du+\xi\left(\int_{U}\pi(u)du-1\right)\right]$
$\qquad=\sup_{\pi\in{\cal P}(U)}\int_{U}\left(q(u)\pi(u)+\xi\pi(u)-\frac{1}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du-\xi$
$\qquad\leq\sup_{\pi(\cdot)>0}\int_{U}\left(q(u)\pi(u)+\xi\pi(u)-\frac{1}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du-\xi$
$\qquad\leq\int_{U}\sup_{\pi(\cdot)>0}\left(q(u)\pi(u)+\xi\pi(u)-\frac{1}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du-\xi.$
Taking $\xi=\psi+\frac{1}{p-1}$ in the previous result, we deduce that the unique maximizer of the inner optimization is given by $\pi^{*}(u)=\left(\frac{p-1}{p\gamma}\right)^{\frac{1}{p-1}}\left(q(u)+\psi\right)_{+}^{\frac{1}{p-1}}$, $\forall\leavevmode\nobreak\ u\in U$. It follows from (2.34) that $\pi^{*}\in{\cal P}(U)$, which implies the uniqueness of the maximizer. ∎
@@ 3 Continuous-time q-Function and Martingale Characterization under Tsallis Entropy
Given $\pi\in\Pi$ and $(t,x,u)\in[0,T]\times\mathds{R}\times U$, let us consider a “perturbed” policy of $\pi$, denoted by $\tilde{\pi}$, as follows: for $\Delta t>0$, it takes the action $u\in U$ on $t,t+\Delta t)$ and then follows $\pi$ on $t+\Delta t,T]$. The resulting state process $X^{\tilde{\pi}}$ with $X_{t}^{\tilde{\pi}}=x$ can be split into two pieces. On $t,t+\Delta t)$, $X^{\tilde{\pi}}=X^{u}$, which is the solution to the following equation: $X^{u}_{t}=x$, and for $s\in[t,t+\Delta t)$,
$dX_{s}^{u}=b(s,X_{s}^{u},u_{s})ds+\sigma(s,X_{s}^{u},u_{s})dW_{s}+{\int_{\mathds{R}}\varphi(s,X_{s-}^{u},u_{s},z)N(ds,dz)},$
@@
The $q$-function of problem (2.9) associated with a given policy $\pi\in\Pi_{t}$ is defined as, for all $(t,x,u)\in[0,T]\times\mathds{R}\times U$,
@@
where the Lagrange multiplier $\psi(t,x):[0,T]\times\mathds{R}\to\mathds{R}$ satisfies
@@
A natural question is whether such a function $\psi(t,x)$ exists. We also emphasize that the function $\psi(t,x)$ can dependent on the policy $\pi$ in view of (3.2) .In fact, given a policy $\pi\in\Pi_{0}$, for fixed $(t,x)\in[0,T]\times\mathds{R}$, let us introduce the following mapping given by
@@
Suppose that Assumptions (A1), (A2) and (A3) hold. Let $\pi\in\Pi_{0}$, its value function $J(\cdot,\cdot;\pi)\in C^{1,2}([0,T)\times\mathds{R})\cap C([0,T]\times\mathds{R})$ and a continuous function $\hat{q}:[0,T]\times\mathds{R}\times U\to\mathds{R}$ be given. Then, we have
@@
    $\hat{q}(t,x,u)=q(t,x,u;\pi)$ for all $(t,x,u)\in[0,T]\times\mathds{R}\times U$ if and only if for all $(t,x)\in[0,T]\times\mathds{R}$ and any time grid $G_{t:T}$ on $[t,T]$, the following process
@@
    If $\hat{q}(t,x,u)=q(t,x,u;\pi)$ for all $(t,x,u)\in[0,T]\times\mathds{R}\times{U}$, given any $\pi^{\prime}\in\Pi_{0}$, for all $(t,x)\in[0,T]\times\mathds{R}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the following process
@@
    If there exists $\pi^{\prime}\in\Pi_{0}$ such that for all $(t,x)\in[0,T]\times\mathds{R}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the process (3.4) is an $(\mathbb{F}^{\prime},\mathbb{Q})$-martingale with initial condition $X_{t}^{\pi^{\prime}}=x$, then $\hat{q}(t,x,u)=q(t,x,u;\pi)$ for all $(t,x,u)\in[0,T]\times\mathds{R}\times U$.
@@
$\int_{U}\left(q(t,x,u;\pi)+\gamma l_{p}(\pi(u|t,x))\right)\pi(u|t,x)du=0,\quad\forall(t,x)\in[0,T]\times\mathds{R}.$ (3.5)
@@
$\quad+\int_{s}^{t}\int_{\mathds{R}}\left(J(l,X_{l-}^{\pi}+\varphi(l,X_{l-}^{\pi},u_{l}^{\pi},z))-J(l,X_{l-}^{\pi})\right)\tilde{N}(dl,dz)+\int_{t}^{s}J_{x}(l,X_{l}^{\pi})\sigma(l,X_{l}^{\pi},u_{l}^{\pi})dW_{l}$
$=$ $\int_{t}^{s}\left[q(l,X_{l}^{\pi},u_{l}^{\pi})-\hat{q}(l,X_{l}^{\pi},u_{l}^{\pi})\right]dl+\int_{t}^{s}J_{x}(l,X_{l}^{\pi})\sigma(l,X_{l}^{\pi},u_{l}^{\pi})dW_{l}$
$\quad+\int_{s}^{t}\int_{\mathds{R}}\left(J(l,X_{l-}^{\pi}+\varphi(l,X_{l-}^{\pi},u_{l}^{\pi},z))-J(l,X_{l-}^{\pi},u_{l}^{\pi})\right)\tilde{N}(dl,dz).$
@@
Define $h(t,x,u)=q(t,x,u;\pi)-\hat{q}(t,x,u)$, a continuous function on $0,T]\times\mathds{R}\times U$. Suppose the conclusion fails: there exist $(t^{*},x^{*},u^{*})\in[0,T)\times\mathds{R}\times U$ and $\epsilon>0$ with $h(t^{*},x^{*},u^{*})>\epsilon$. By continuity of $h$, there exists $\delta>0$ such that $h(l,x^{\prime},u^{\prime})>\epsilon/2$ whenever $|l-t^{*}|\vee|x^{\prime}-x^{*}|\vee|u^{\prime}-u^{*}|<\delta$ (where $u\vee v=\max\{u,v\}$).
@@
$\tau=\inf\left\{l\geq t^{*};\leavevmode\nobreak\ |l-t^{*}|>\delta\text{ or }|X_{l}^{\pi}-x^{*}|>\delta\right\}=\inf\left\{l\geq t^{*};\leavevmode\nobreak\ |X_{l}^{\pi}-x^{*}|>\delta\right\}\wedge(t^{*}+\delta),$
with $u\wedge v=\min\{u,v\}$. By continuity of $X^{\pi}$, $\tau>t^{*}$, $\mathbb{Q}$-a.s.
@@
contradicting (3). This proves that $q(t,x,u;\pi)=\hat{q}(t,x,u)$ for every $(t,x,u)\in[0,T]\times\mathds{R}\times U$.
The proofs of (ii) and (iii) are parallel to the first part of the proof of (i); thus, we omit them here. Moreover, Eq. (3.5) follows from the definition of the q-function in Definition 3.1 and the PDE (2.3). ∎
@@
Let Assumptions (A1), (A2) and (A3) hold. For each $p\geq 1$, let a policy $\pi_{p}\in\Pi_{0}$, a function $\hat{J}\in C^{1,2}([0,T)\times\mathds{R})\cap C([0,T]\times\mathds{R})$ and a continuous function $\hat{q}:[0,T]\times\mathds{R}\times U\to\mathds{R}$ be given such that, for all $(t,x)\in[0,T]\times\mathds{R}$,
@@
    $\hat{J}$ and $\hat{q}$ are respectively the value function satisfying (2.3) and the $q$-function associated with $\pi_{p}$ if and only if for all $(t,x)\in[0,T]\times\mathds{R}$ and any time grid $G_{t:T}$ on $[t,T]$, the following process
@@
    If $\hat{J}$ and $\hat{q}$ are respectively the value function and the $q$-function associated with $\pi_{p}$, given any $\pi^{\prime}\in\Pi_{0}$, for all $(t,x)\in[0,T]\times\mathbb{R}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the following process
@@
    If there exists $\pi^{\prime}\in\Pi_{0}$ such that, for all $(t,x)\in[0,T]\times\mathbb{R}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the process (3.8) is an $(\mathbb{F},\mathbb{Q})$-martingale, where $(X_{s}^{\pi^{\prime}})_{s\in[t,T]}$ is the solution to (2.1) under $\pi^{\prime}$ with $X_{t}^{\pi^{\prime}}=x$, then $\hat{J}$ and $\hat{q}$ are respectively the value function and the $q$-function associated with $\pi_{p}$.
@@
with the normalizing function $\psi(t,x)$ satisfying $\int_{U}\left(\frac{p-1}{p\gamma}\right)^{\frac{1}{p-1}}\left(\hat{q}(t,x,u)+\psi(t,x)\right)_{+}^{\frac{1}{p-1}}du=1$ for all $(t,x)\in[0,T]\times\mathds{R}$, then $\pi_{p}$ ($p\geq 1$) is an optimal policy and $\hat{J}$ is the corresponding optimal value function.
@@
$=$ $\int_{t}^{s}\Big{[}\hat{J}_{t}(l,X_{l}^{\pi_{p}})dl+{\cal H}(l,X_{l}^{\pi_{p}},u_{l}^{\pi_{p}},\hat{J}(l,X_{l}^{\pi_{p}}),\hat{J}_{x}(l,X_{l}^{\pi_{p}}),\hat{J}_{xx}(l,X_{l}^{\pi_{p}}))$
$\quad-\hat{q}(l,X_{l}^{\pi_{p}},{\pi_{p}})\Big{]}dl+\int_{s}^{t}\int_{\mathds{R}}\left(\hat{J}(l,X_{l-}^{\pi_{p}}+\varphi(l,X_{l-}^{\pi_{p}},u_{l}^{\pi_{p}},z))-\hat{J}(l,X_{l-}^{\pi_{p}})\right)\tilde{N}(dl,dz)$
$\quad+\int_{t}^{s}\hat{J}_{x}(l,X_{l}^{\pi_{p}})\sigma(l,X_{l}^{\pi_{p}},u_{l}^{\pi_{p}})dW_{l}$
$=-\int_{t}^{s}\left[\hat{q}(l,X_{l}^{\pi_{p}};{\pi_{p}})+\gamma l_{p}({\pi_{p}})\right]dl+\int_{t}^{s}\hat{J}_{x}(l,X_{l}^{\pi_{p}})\sigma(l,X_{l}^{\pi_{p}},u_{l}^{\pi_{p}})dW_{l}$
$\quad+\int_{s}^{t}\int_{\mathds{R}}\left(\hat{J}(l,X_{l-}^{\pi_{p}}+\varphi(l,X_{l-}^{\pi_{p}},u_{l}^{\pi_{p}},z))-\hat{J}(l,X_{l-}^{\pi_{p}})\right)\tilde{N}(dl,dz).$
@@
We now prove the other side. Assume that (3.7) is an $(\mathbb{F},\mathbb{Q})$-martingale. Using Proposition 3.2-(i), we derive that $\hat{q}(t,x,u)=\hat{J}_{t}(t,x)+{\cal H}(t,x,u,\hat{J},\hat{J}_{x},\hat{J}_{xx})$, for $(t,x,u)\in[0,T]\times\mathds{R}\times U$. Then, the constraint (3.6) reads
@@
1:Initialize $\theta$$\zeta$ and $i=1$.
@@
In this subsection, we handle the case when the normalizing function $\psi(t,x)$ does not admit an explicit form. In this case, by knowing the learned q-function, we cannot learn the optimal policy directly as there is an unknown term $\psi(t,x)$. We can still parameterize the value function by a family of functions $J^{\theta}(\cdot)$, where $\theta\in\Theta\subset\mathds{R}^{L_{\theta}}$ and $L_{\xi}$ is the dimension of the parameter, and parameterize the q-function by a family of functions $q^{\zeta}(\cdot,\cdot)$, where $\zeta\in\Psi\subset\mathds{R}^{L_{\zeta}}$ and $L_{\zeta}$ is the dimension of the parameter. However, we can not get the normalizing function $\psi(t,x)$ from (4.1). In response, we parameterize the policy by a family policy function $\pi^{\chi}(\cdot)$, where $\chi\in\Upsilon\subset\mathds{R}^{L_{\chi}}$ and $L_{\chi}$ is the dimension of the parameter. Moreover, the approximators $J^{\theta}$ and $\pi^{\chi}$ should also satisfy $J^{\theta}(T,x)=g(x)$. Define the function $F:[0,T]\times\mathds{R}\times{\cal P}(U)\times{\cal P}(U)\to\mathds{R}$ by

### v4

@@ 2.3 Exploratory HJB equation and policy improvement iteration
Let $\xi=\psi-\frac{\gamma}{p-1}$, and consider the following optimization problem
$\sup_{\pi\in{\cal P}(U)}\left[\int_{U}\left(q(u)\pi(u)+\frac{\gamma}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du+\xi\left(\int_{U}\pi(u)du-1\right)\right].$ (2.37)
As $\int_{U}\pi(u)du=1$ for $\pi\in{\cal P}(U)$, the optimization problems (2.36) and (2.37) are equivalent. For problem (2.37), we have
$\sup_{\pi\in{\cal P}(U)}\left[\int_{U}\left(q(u)\pi(u)+\frac{\gamma}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du+\xi\left(\int_{U}\pi(u)du-1\right)\right]$
$\qquad=\sup_{\pi\in{\cal P}(U)}\int_{U}\left(q(u)\pi(u)+\xi\pi(u)+\frac{\gamma}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du-\xi$
$\qquad\leq\sup_{\pi(\cdot\geq 0}\int_{U}\left(q(u)\pi(u)+\xi\pi(u)+\frac{\gamma}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du-\xi$
$\qquad\leq\int_{U}\sup_{\pi(\cdot)\geq 0}\left(q(u)\pi(u)+\xi\pi(u)+\frac{\gamma}{p-1}\left(\pi(u)-\pi^{p}(u)\right)\right)du-\xi.$ (2.38)
We can then deduce that the unique maximizer of the optimization problem on the right-hand side of (2.3) is given by
$\pi^{*}(u)=\left(\frac{p-1}{p\gamma}\right)^{\frac{1}{p-1}}\left(q(u)+\psi\right)_{+}^{\frac{1}{p-1}},\quad\text{for}\penalty 10000\ u\in U.$
It follows from (2.35) that $\pi^{*}\in{\cal P}(U)$, which completes the proof. ∎
@@ 3 Continuous-time q-Function and Martingale Characterization under Tsallis Entropy
Given $\pi\in\Pi$ and $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U$, let us consider a “perturbed” policy of $\pi$, denoted by $\tilde{\pi}$, as follows: for $\Delta t>0$, it takes the action $u\in U$ on $t,t+\Delta t)$ and then follows $\pi$ on $t+\Delta t,T]$. The resulting state process $X^{\tilde{\pi}}$ with $X_{t}^{\tilde{\pi}}=x$ can be split into two pieces. On $t,t+\Delta t)$, $X^{\tilde{\pi}}=X^{u}$, which is the solution to the following equation: $X^{u}_{t}=x$, and for $s\in[t,t+\Delta t)$,
$dX_{s}^{u}=b(s,X_{s}^{u},u_{s})ds+\sigma(s,X_{s}^{u},u_{s})dW_{s}+{\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\varphi(s,X_{s-}^{u},u_{s},z){\bf 1}_{\{r\leq\lambda(s,X_{s-}^{u},u_{s}\}}N(ds,dz,dr)},$
@@
The $q$-function of problem (2.9) associated with a given policy $\pi\in\Pi_{t}$ is defined as, for all $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U$,
@@
where the Lagrange multiplier $\psi(t,x):[0,T]\times\mathds{R}^{n}\to\mathds{R}$ satisfies
@@
A natural question is whether such a function $\psi(t,x)$ exists. We also emphasize that the function $\psi(t,x)$ can dependent on the policy $\pi$ in view of (3.2). In fact, given a policy $\pi\in\Pi_{0}$, for fixed $(t,x)\in[0,T]\times\mathds{R}^{n}$, let us introduce the following mapping given by
@@
Suppose that Assumptions (A1), (A2) and (A3) hold. For a policy $\pi\in\Pi_{0}$, assume that its value function $J(\cdot,\cdot;\pi)\in C^{1,2}([0,T)\times\mathds{R}^{n})\cap C([0,T]\times\mathds{R}^{n})$. Given a continuous function $\hat{q}:[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}$. Then, we have
@@
    $\hat{q}(t,x,u)=q(t,x,u;\pi)$ for all $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U$ if and only if for all $(t,x)\in[0,T]\times\mathds{R}^{n}$ and any time grid $G_{t:T}$ on $[t,T]$, the following process
@@
    If $\hat{q}(t,x,u)=q(t,x,u;\pi)$ for all $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times{U}$, given any $\pi^{\prime}\in\Pi_{0}$, for all $(t,x)\in[0,T]\times\mathds{R}^{n}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the following process
@@
    If there exists $\pi^{\prime}\in\Pi_{0}$ such that for all $(t,x)\in[0,T]\times\mathds{R}^{n}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the process (3.4) is an $(\mathbb{F}^{\prime},\mathbb{Q})$-martingale with initial condition $X_{t}^{\pi^{\prime}}=x$, then $\hat{q}(t,x,u)=q(t,x,u;\pi)$ for all $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U$.
@@
$\int_{U}\left(q(t,x,u;\pi)+\gamma l_{p}(\pi(u|t,x))\right)\pi(u|t,x)du=0,\quad\forall(t,x)\in[0,T]\times\mathds{R}^{n}.$ (3.5)
@@
$\quad+\int_{t}^{s}\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\left(J(l,X_{l-}^{\pi}+\varphi(l,X_{l-}^{\pi},u_{l}^{\pi},z))-J(l,X_{l-}^{\pi})\right){\bf 1}_{\{r\leq\lambda(l,X_{l-}^{\pi},u_{l}^{\pi})\}}\tilde{N}(dl,dz,dr)$
$\quad+\int_{t}^{s}J_{x}(l,X_{l}^{\pi})^{\top}\sigma(l,X_{l}^{\pi},u_{l}^{\pi})dW_{l}$
$=$ $\int_{t}^{s}\left[q(l,X_{l}^{\pi},u_{l}^{\pi})-\hat{q}(l,X_{l}^{\pi},u_{l}^{\pi})\right]dl+\int_{t}^{s}J_{x}(l,X_{l}^{\pi})^{\top}\sigma(l,X_{l}^{\pi},u_{l}^{\pi})dW_{l}$
$\quad+\int_{t}^{s}\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\left(J(l,X_{l-}^{\pi}+\varphi(l,X_{l-}^{\pi},u_{l}^{\pi},z))-J(l,X_{l-}^{\pi})\right){\bf 1}_{\{r\leq\lambda(l,X_{l-}^{\pi},u_{l}^{\pi})\}}\tilde{N}(dl,dz,dr).$
@@
Define $h(t,x,u)=q(t,x,u;\pi)-\hat{q}(t,x,u)$, a continuous function on $0,T]\times\mathds{R}^{n}\times U$. Suppose the conclusion fails: there exist $(t^{*},x^{*},u^{*})\in[0,T)\times\mathds{R}^{n}\times U$ and $\epsilon>0$ with $h(t^{*},x^{*},u^{*})>\epsilon$. By continuity of $h$, there exists $\delta>0$ such that $h(l,x^{\prime},u^{\prime})>\epsilon/2$ whenever $|l-t^{*}|\vee|x^{\prime}-x^{*}|\vee|u^{\prime}-u^{*}|<\delta$ (where $u\vee v=\max\{u,v\}$).
@@
$\tau=\inf\left\{l\geq t^{*};\penalty 10000\ |l-t^{*}|>\delta\text{ or }|X_{l}^{\pi}-x^{*}|>\delta\right\}=\inf\left\{l\geq t^{*};\penalty 10000\ |X_{l}^{\pi}-x^{*}|>\delta\right\}\wedge(t^{*}+\delta),$
with $u\wedge v=\min\{u,v\}$. By the stochastic continuity of $X^{\pi}$, $\tau>t^{*}$, $\mathbb{Q}$-a.s.
@@
contradicting (3). This proves that $q(t,x,u;\pi)=\hat{q}(t,x,u)$ for every $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U$.
The proofs of (ii) and (iii) are similar to the first part of the proof of (i) and are hence omitted. Moreover, Eq. (3.5) follows from the definition of the q-function in Definition 3.1 and the PDE (2.3). ∎
@@
Let Assumptions (A1), (A2) and (A3) hold. For each $p\geq 1$, let a policy $\pi_{p}\in\Pi_{0}$, a function $\hat{J}\in C^{1,2}([0,T)\times\mathds{R}^{n})\cap C([0,T]\times\mathds{R}^{n})$ and a continuous function $\hat{q}:[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}$ be given such that, for all $(t,x)\in[0,T]\times\mathds{R}^{n}$,
@@
    $\hat{J}$ and $\hat{q}$ are respectively the value function satisfying (2.3) and the $q$-function associated with $\pi_{p}$ if and only if for all $(t,x)\in[0,T]\times\mathds{R}^{n}$ and any time grid $G_{t:T}$ on $[t,T]$, the following process
@@
    If $\hat{J}$ and $\hat{q}$ are respectively the value function and the $q$-function associated with $\pi_{p}$, given any $\pi^{\prime}\in\Pi_{0}$, for all $(t,x)\in[0,T]\times\mathbb{R}^{n}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the following process
@@
    If there exists $\pi^{\prime}\in\Pi_{0}$ such that, for all $(t,x)\in[0,T]\times\mathbb{R}^{n}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the process (3.8) is an $(\mathbb{F},\mathbb{Q})$-martingale, where $(X_{s}^{\pi^{\prime}})_{s\in[t,T]}$ is the solution to (2.1) under $\pi^{\prime}$ with $X_{t}^{\pi^{\prime}}=x$, then $\hat{J}$ and $\hat{q}$ are respectively the value function and the $q$-function associated with $\pi_{p}$.
@@
with the normalizing function $\psi(t,x)$ satisfying $\int_{U}\left(\frac{p-1}{p\gamma}\right)^{\frac{1}{p-1}}\left(\hat{q}(t,x,u)+\psi(t,x)\right)_{+}^{\frac{1}{p-1}}du=1$ for all $(t,x)\in[0,T]\times\mathds{R}^{n}$, then $\pi_{p}$ ($p\geq 1$) is an optimal policy and $\hat{J}$ is the corresponding optimal value function.
@@
$=$ $\int_{t}^{s}\Big[\hat{J}_{t}(l,X_{l}^{\pi_{p}})dl+{\cal H}(l,X_{l}^{\pi_{p}},u_{l}^{\pi_{p}},\hat{J}(l,X_{l}^{\pi_{p}}),\hat{J}_{x}(l,X_{l}^{\pi_{p}}),\hat{J}_{xx}(l,X_{l}^{\pi_{p}}))-\hat{q}(l,X_{l}^{\pi_{p}},{\pi_{p}})\Big]dl$
$\quad+\int_{t}^{s}\int_{\mathds{R}}\left(\hat{J}(l,X_{l-}^{\pi_{p}}+\varphi(l,X_{l-}^{\pi_{p}},u_{l}^{\pi_{p}},z))-\hat{J}(l,X_{l-}^{\pi_{p}})\right){\bf 1}_{\{r\leq\lambda(l,X_{l-}^{\pi_{p}},u_{l}^{\pi_{p}})\}}\tilde{N}(dl,dz,dr)$
$\quad+\int_{t}^{s}\hat{J}_{x}(l,X_{l}^{\pi_{p}})^{\top}\sigma(l,X_{l}^{\pi_{p}},u_{l}^{\pi_{p}})dW_{l}$
$=-\int_{t}^{s}\left[\hat{q}(l,X_{l}^{\pi_{p}};{\pi_{p}})+\gamma l_{p}({\pi_{p}})\right]dl+\int_{t}^{s}\hat{J}_{x}(l,X_{l}^{\pi_{p}})^{\top}\sigma(l,X_{l}^{\pi_{p}},u_{l}^{\pi_{p}})dW_{l}$
$\quad+\int_{t}^{s}\int_{\mathds{R}}\left(\hat{J}(l,X_{l-}^{\pi_{p}}+\varphi(l,X_{l-}^{\pi_{p}},u_{l}^{\pi_{p}},z))-\hat{J}(l,X_{l-}^{\pi_{p}})\right){\bf 1}_{\{r\leq\lambda(l,X_{l-}^{\pi_{p}},u_{l}^{\pi_{p}})\}}\tilde{N}(dl,dz,dr).$
@@
We now prove the other side. Assume that (3.7) is an $(\mathbb{F},\mathbb{Q})$-martingale. Using Proposition 3.2-(i), we derive that $\hat{q}(t,x,u)=\hat{J}_{t}(t,x)+{\cal H}(t,x,u,\hat{J},\hat{J}_{x},\hat{J}_{xx})$, for $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U$. Then, the constraint (3.6) reads
@@
Remark 3.4.
Proposition 3.2 and Theorem 3.3 show that the form of the martingale conditions are similar to those under Shannon entropy, which are robust to the specific choice of entropy.
Next, we present the martingale characterization of the optimal value function $J^{*}$ and the optimal q-function. The proof of Theorem 3.5 closely follows those of Proposition 3.2 and Theorem 3.3, and is hence omitted.
Theorem 3.5.
Let Assumptions (A1), (A2) and (A3) hold. For each $p\geq 1$, given a policy $\pi_{p}\in\Pi_{0}$, a function $\hat{J}\in C^{1,2}([0,T)\times\mathds{R}^{n})\cap C([0,T]\times\mathds{R}^{n})$ and a continuous function $\hat{q}:[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}$ such that, for all $(t,x)\in[0,T]\times\mathds{R}^{n}$,
$\hat{J}(T,x)=g(x),$ (3.10)
$\left(\frac{p-1}{p\gamma}\right)^{\frac{p}{p-1}}\int_{U}\left(\left(\hat{q}(t,x,u)+\psi(t,x)\right)_{+}-p\hat{q}(t,x,u)\right)\left(\hat{q}(t,x,u)+\psi(t,x)\right)_{+}^{\frac{1}{p-1}}du=1,$ (3.11)
where $\psi(t,x)$ is the normalizing function satisfying $\int_{U}\left(\frac{p-1}{p\gamma}\right)^{\frac{1}{p-1}}\left(\hat{q}(t,x,u)+\psi(t,x)\right)_{+}^{\frac{1}{p-1}}du=1$ for all $(t,x)\in[0,T]\times\mathds{R}^{n}$. Then, it holds that
  • (i)
    If $\hat{J}^{*}$ and $\hat{q}^{*}$ are respectively the optimal value function and the optimal $q$-function, given any $\pi^{\prime}\in\Pi_{0}$, for all $(t,x)\in[0,T]\times\mathbb{R}^{n}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the following process
    $\hat{J}^{*}\left(s,X_{s}^{\pi^{\prime}}\right)+\int_{t}^{s}\left[f\left(l,X_{l}^{\pi^{\prime}},u_{l}^{\pi^{\prime}}\right)-\hat{q}^{*}\left(l,X_{l}^{\pi^{\prime}},u_{l}^{\pi^{\prime}}\right)\right]dl,\quad s\in[t,T]$ (3.12)
    is an $(\mathbb{F},\mathbb{Q})$-martingale. Here, $X^{\pi^{\prime}}=(X_{s}^{\pi^{\prime}})_{s\in[t,T]}$ is the solution to (2.1) under $\pi^{\prime}$ with $X_{t}^{\pi^{\prime}}=x$. Furthermore, the policy $\pi^{*}_{p}$ given by
    $\pi_{p}^{*}(u|t,x)=\left(\frac{p-1}{p\gamma}\right)^{\frac{1}{p-1}}\left(\hat{q}(t,x,u)+\psi(t,x)\right)_{+}^{\frac{1}{p-1}},\quad p\geq 1$ (3.13)
    is the optimal policy.
  • (ii)
    If there exists some $\pi^{\prime}\in\Pi_{0}$ such that, for all $(t,x)\in[0,T]\times\mathbb{R}^{n}$ and any time grid ${G}_{t:T}$ on $[t,T]$, the process (3.12) is an $(\mathbb{F},\mathbb{Q})$-martingale, where $(X_{s}^{\pi^{\prime}})_{s\in[t,T]}$ is the solution to (2.1) under $\pi^{\prime}$ with $X_{t}^{\pi^{\prime}}=x$, then $\hat{J}^{*}$ and $\hat{q}^{*}$ are respectively the optimal value function and the optimal $q$-function.
@@
1:Initialize $\theta$, $\zeta$ and $i=1$.
@@
In this subsection, we handle the case when the normalizing function $\psi(t,x)$ does not admit an explicit form. In this case, by knowing the learned q-function, we cannot learn the optimal policy directly as there is an unknown term $\psi(t,x)$. We can still parameterize the value function by a family of functions $J^{\theta}(\cdot)$, where $\theta\in\Theta\subset\mathds{R}^{L_{\theta}}$ and $L_{\xi}$ is the dimension of the parameter, and parameterize the q-function by a family of functions $q^{\zeta}(\cdot,\cdot)$, where $\zeta\in\Psi\subset\mathds{R}^{L_{\zeta}}$ and $L_{\zeta}$ is the dimension of the parameter. However, we can not get the normalizing function $\psi(t,x)$ from (4.1). In response, we parameterize the policy by a family policy function $\pi^{\chi}(\cdot)$, where $\chi\in\Upsilon\subset\mathds{R}^{L_{\chi}}$ and $L_{\chi}$ is the dimension of the parameter. Moreover, the approximators $J^{\theta}$ and $\pi^{\chi}$ should also satisfy $J^{\theta}(T,x)=g(x)$. Define the function $F:[0,T]\times\mathds{R}^{n}\times{\cal P}(U)\times{\cal P}(U)\to\mathds{R}$ by